A shirt was originally priced at $30. The store is having a 15% off sale. How much will you pay for the shirt after the discount (not including tax)? Round your answer to the nearest cent
A $30 shirt marked 15% off will cost

Answer by ikleyn(52748)

.
A shirt was originally priced at $30. The store is having a 15% off sale.
How much will you pay for the shirt after the discount (not including tax)?
Round your answer to the nearest cent
A $30 shirt marked 15% off will cost
~~~~~~~~~~~~~~~~

It will cost 30 - 0.15*30 = (1-0.15)*30 = 0.85*30 = 25.50 dollars, not including tax.    ANSWER

Solved.